NatalieAfton:

Examine the system of equations. y = 3 2 x − 6, y = −9 2 x + 21 Use substitution to solve the system of equations. What is the value of y? y =

yourgrandpa:

By substitution, 1.5x - 6 = -4.5x + 21 6x = 27 x = 4.5 Hence, y = 1.5(4.5) - 6 = 0.75

NatalieAfton:

thx it's also 3/4

yourgrandpa:

yes (:

yourgrandpa:

The correct answer is 3/4 you were right